An excited hydrogen atom releases an electromagnetic wave to return to its normal state. You use your futuristic dual electric/magnetic field tester on the electromagnetic wave to find the directions of the electric field and magnetic field. Your device tells you that the electric field is pointing in the positive xdirection and the magnetic field is pointing in the positive z direction. In which direction does the released electromagnetic wave travel?
Answer #1

In an electromagnetic wave the direction of electric field, the direction of magnetic field and the direction of propagation are mutually perpendicular.

The direction of propagation of wave is the direction of \bold{E}\times\bold{B}

\hat{i},\hat{j},\hat{k} are the unit vectors along positive X-axis , positive Y-axis , positive Z-axis respectively.

Given that  \bold{E}=E\hat{i} and \bold{B}=B\hat{k}

The direction of propagation of electromagnetic wave is \bold{E}\times\bold{B}=E\hat{i}\times B\hat{k}=EB(-\hat{j})

That is electromagnetic wave travels along negative Y-axis.